# Induced Berry Connection and Photonic Spin Hall Effect in Optical Dirac Theory

### v1

@@ I Introduction
Polarization refers to the orientation of an electromagnetic wave’s electric/magnetic field vector. In the context of quantum theory, it corresponds to the spin angular momentum of photons, an intrinsic degree of freedom of light encoded in the classical Maxwell theory. Adapting our understanding of electromagnetic waves to incorporate the concept of angular momentum is critical for comprehending the intricate dynamics at play. Besides the SAM of an electromagnetic wave packet, we have appreciated another two pivotal components: intrinsic orbital angular momentum associated with optical vortices, characterizes the inherent rotation of light around its axis, and extrinsic orbital angular momentum corresponds to the helical optical trajectory, extending beyond its inherent properties to encompass its path through space [1, 2, 3, 4, 5, 6, 7]. Interactions among these three angular momenta give rise to a diverse range of optical phenomena and can be manipulated to control various aspects of optical systems.
@@
The spin-orbit interaction (SOI) of light manifests itself through the interplay and mutual conversion between three types of OAM [8, 9, 10, 11, 12, 13], a family of fundamental phenomena that arises directly from the solutions of Maxwell’s equations, particularly when dealing with wavelength-scale micro- and nanostructures. Obviously, a full treatment of the SOI phenomenon on these scales requires vector-wave mechanics. This motivation has led us to develop an extended Dirac theory for optical fields in generic media, which was found to be a non-Hermitian chiral extension of massive fermions with anomalous magnetic moments moving in an external pseudo-magnetic field [14]. In this approach, an optical medium prescribed by a symmetric dielectric tensor can be decomposed in terms of the monopole, dipole, and quadruple components, corresponding to spin-0 scalar mode, spin-1 vector mode, and spin-2 tensor mode, respectively. While coupling to optical fields, those virtual excitation modes ensure the conservation of total angular momentum (TAM). On the other hand, the chiral $\gamma_{5}$ extension leads to a scenario of an anomalous magnetic momentum in an equivalent magnetic field, a synthetic one in a non-spin-degenerate medium.
@@
This study aims to develop a field-theoretical representation of spin-orbit interaction (SOI) and the spin/orbit Hall effects (SHE/OHE) within the framework of optical Dirac theory, providing deeper physical insights into related optical phenomena. To this end, we extend optical Dirac theory [14] by incorporating antisymmetric components in the dielectric tensor. These components manifest either as a rotational vector in the spin-1 mode, associated with a helical optical path, or as a spin-2 tensor mode, corresponding to gyro-electric/magnetic media. In Section II, we derive the optical Dirac equation in a helical coordinate system with a rotational dipole-vector, leading to an effective Hamiltonian that explicitly captures both SAM-EOAM and IOAM-EOAM interactions. We also provide leading-order solutions to the optical Dirac equation, incorporating SAM/IOAM-EOAM interactions. By retaining longitudinal corrections, we further examine back-scattering phenomena [15] related to the anti-photon state. In Section III, we employ the Heisenberg equation to derive the photon motion equation, introducing characteristic corrections to light-ray trajectories, which manifest as the spin Hall effect (SHE) and orbital Hall effect (OHE). Section IV explores the evolution of SAM and IOAM in the Heisenberg picture, elucidating their role in the broader SHE/OHE family. In Section V, we analyze the interaction between helicity-independent spin and EOAM, revealing behavior analogous to that of IOAM. Finally, Section VI summarizes our findings and presents concluding remarks. The extension to gyro-electric/magnetic media follows naturally, as shown in Appendix A.
@@ II OPTICAL DIRAC EQUATION IN HELICAL COORDINATES
For light propagation in a gravitational field, the curved space background can be transformed to a linear optical medium, whose optical properties characterized by the effective dielectric tensor are fully specified by the spacetime geometries [16]. Typically, in a rotating gravitational field, a well-known optical phenomenon is the classical phase shift referred to as the Sagnac factor [17], currently explained by the dragging effect of the frame in general relativity [18], which can be understood alternatively by the coupling between the rotation and the OAM of photons $\sim{\bf\Omega}\cdot{\bf L}$. In addition, another phenomenon is an extra polarization-dependent phase shift resulting from the vortical dragging vector, which leads to the optical activity of rotating gravitational fields [19, 20, 21]. Basically, it can also be equivalently described by the coupling between the macroscopic rotation ${\bf\Omega}$ and the internal spin ${\bf s}$ of the photon [22, 20]. This coupling has a typical form of $\sim{\bf\Omega}\cdot{\bf s}$.
Figure 1: The local coordinate system is induced from the Frenet frame $\mathbf{t},\mathbf{n},\mathbf{b}$. Here, $\mathbf{r}_{c}$ denotes the trajectory of the center of mass, and $\mathbf{r}_{\perp}$ is a transverse displacement vector relative to the center of mass.
It is noted that both rotation-OAM/SAM couplings are due to spatial rotation in time, which can be attributed to the dragging vector, the space-time component ${\bf g}=\{g_{0i}\}$ in metric. Similarly, a similar rotation occurs in space for a helical optical path. As demonstrated later, this rotation around a spatial axis will give rise to a similar coupling effect, e.g., the optical SHE. Suppose light propagates along the helical path in space (see FIG. 1). Under the geometric optics approximation, an optical path can be parameterized by its optical length $s$, and let $\mathbf{r}_{\perp}=\{x_{1},x_{2}\}$ denote a transverse displacement vector around the propagating direction, we can introduce local coordinates $(s,x_{1},x_{2})$ such that
@@
In the local Frenet-Serret orthonormal tetrad defined by curvature $\kappa(s)$ and torsion $\tau(s)$, the line element is $d{\bf R}^{2}=\gamma_{\alpha\beta}dx^{\alpha}dx^{\beta}$ where Greek indices run over 0, 1 and 2, and $x_{0}=s$ by convention. The metric components can be easily found [23]:
$\gamma_{00}=(1-\kappa x_{1})^{2}+\tau^{2}r_{\perp}^{2},\gamma_{0\alpha}=-\tau\varepsilon_{0\alpha\beta}x_{\beta},\gamma_{\alpha\beta}=\delta_{\alpha\beta}$ (2)
@@
${\frac{1}{\sqrt{\gamma}}\frac{\partial}{\partial t}}{(\sqrt{\gamma}D^{\alpha})}=\epsilon^{\alpha\beta\rho}\nabla_{\beta}\bigr{(}\sqrt{\gamma}\gamma_{\rho\sigma}B^{\sigma}\bigr{)}$ (3)
${\frac{1}{\sqrt{\gamma}}\frac{\partial}{\partial t}}{(\sqrt{\gamma}B^{\alpha})}=\epsilon^{\alpha\beta\rho}\nabla_{\beta}\bigr{(}\sqrt{\gamma}\gamma_{\rho\sigma}D^{\sigma}\bigr{)}$
${\frac{1}{\sqrt{\gamma}}}\nabla_{\alpha}(\sqrt{\gamma}D^{\alpha})=0$
${\frac{1}{\sqrt{\gamma}}}\nabla_{\alpha}(\sqrt{\gamma}B^{\alpha})=0$
where $\sqrt{\gamma}=|\det(\gamma_{\alpha\beta})|^{1/2}$ that can be absorbed into the definition of the electromagnetic vectors ${\bf D}$ and ${\bf B}$ by this rescaling. It is noted that an electromagnetic system described by Eq.(3) in curvilinear coordinates is equivalent to an optical medium with identical permittivity and permeability tensors,
$\epsilon^{-1}_{\alpha\beta}=\mu^{-1}_{\alpha\beta}=\gamma_{\alpha\beta}$ (4)
which is currently referred to as the spin-degenerate condition in photonic topological insulators. Following the same method as described in Feng & Wu [14], we project the Maxwell equation onto the helicity space spanned by ${\bf e}_{\pm}=({\bf e}_{1}\pm i{\bf e}_{2})/\sqrt{2}$ where $\{{\bf e}_{1},{\bf e}_{2}\}$ is given by Eq.(1) to define the transverse tetrad perpendicular to the wavenumber vector of photon ${\bf e}_{\bf k}={\bm{\nabla}}_{s}$. Let $\{\bm{D}_{\perp},{\bm{B}}_{\perp}\}$ denote the transverse components of electromagnetic field vectors in the helicity space, we introduce two-component wavefunctions for photons in the transverse helicity space:
@@
The derivation of the optical Dirac equation in a generic medium with antisymmetric components, along with its application to the helical coordinate system Eq.(6), is thoroughly detailed in Appendix A. In the curvilinear coordinate specified by Eq.(2), we have
@@
in which the first term in $\gamma_{0}$ term of r.h.s. gives the leading order of the Hamiltonian, including the usual propagating wavenumber $\hat{k}_{z}$ and the rotation-angular momentum coupling, ${\mathbf{\Omega}}\cdot{\mathbf{J}}$, where the rotation is specified by the torsion of helical path, $\mathbf{\Omega}=-\tau{\bf e}_{z}$, and ${\bf J}$ is the TAM, ${\bf J}={\bf L}+{\bf\Sigma}$ and ${\bf\Sigma}=\bm{\sigma}\otimes\sigma_{0}$; The second term $\propto 1/k_{z}$ results from the paraxial corrections, $P_{0}=P_{I}+P_{3}\sigma_{3}$, $P_{\perp}\cdot\sigma_{\perp}=P_{+}\sigma_{+}+P_{-}\sigma_{-}$. The explicit forms of $P_{0},P_{\perp},P_{3}$ in the helical coordinate are given in Eq.(118)-(120) of Appendix A.
@@
$\mathbf{k}_{\perp}=i\kappa_{x}\mathbf{e}_{x}+i\kappa_{y}\mathbf{e}_{y}.$ (8)
According to the Hamiltonian read from Eq.(6), a non-Hermitian scalar term $-i{\bf g}_{\perp}/k_{z}$ appears in $P_{0}$ (see Eq.(118)). However, when coupled to the evanescent wave, this term transforms into a Hermitian wave-interaction ${\bm{\kappa}}_{g}\cdot{\bm{\kappa}}$ where ${\bm{\kappa}}_{g}={\bf g}_{\perp}/k_{z}$, and contributes an additional energy correction.
@@
with the 2-spinors $\chi_{1}=\displaystyle{\binom{1}{0}}$ and $\chi_{2}=\displaystyle{\binom{0}{1}}$. Keeping the leading order of $O(k_{z})$, we can easily see that the eigenvalues for the positive/negative energy states $u_{\alpha}$ and $v_{\alpha}$ are $\omega_{\alpha}^{\pm}=\pm k$, respectively. In Dirac’s relativistic theory of the electron, the negative energy state corresponds to its anti-particle, with opposite charge, namely, the positron. However, for the spin-1 photon, the corresponding antiparticle is itself. Obviously, it can be understood as moving toward the negative arrow of time, equivalently, a backward propagation solution.
@@
$\omega^{(+)}=\omega_{\alpha}^{+}=k_{z}\mp\tau$ (11)
@@
$\omega^{(-)}=-\omega_{\alpha}^{-}=-(k_{z}\pm\tau)$ (12)
@@
$\omega_{\alpha}^{+}$ $=k_{z}\mp(l\tau+\tau)=k_{z}\mp(l+1)\tau$ (14)
$\omega_{\alpha}^{-}$ $=k_{z}\mp(l\tau-\tau)=k_{z}\mp(l-1)\tau$
The first two equations correspond to the left (right) spin of the normal state photon, and the last two equations correspond to the left (right) spin of the negative state photon. We focus exclusively on vortex solutions:
$|\Psi\rangle=e^{-i\omega t+ikz}e^{il\varphi}(\psi_{R}|e_{+}\rangle+\psi_{L}|e_{-}\rangle)$ (15)
the eigen-energies, incorporating AM-rotation coupling $\mathbf{\Omega}\cdot{\bf J}$, are given by:
$\omega=k-j_{z}\tau$ (16)
with $j_{z}$ representing the component of the TAM $\mathbf{J}=\mathbf{L}+\mathbf{\Sigma}$ along the propagate direction. The possible values of $j_{z}$ are $j_{z}\in\{-l-1,-l+1,l-1,l+1\}$. The second term contributes to Berry phase as [25]:
@@
$ik_{z}\frac{\partial\xi}{\partial z}=-\frac{1}{2}g_{s}\nabla_{\perp}^{2}\xi-\hat{h}_{+}\xi-\mathbf{P}_{\perp}\cdot\mathbf{\sigma}_{\perp}\xi$ (22)
in which $\mathbf{P}_{\perp}\cdot\mathbf{\sigma}_{\perp}=P_{+}\sigma_{+}+P_{-}\sigma_{-}$. Similarly, defining
@@
$ik_{z}\frac{\partial\eta}{\partial z}=-\frac{1}{2}g_{s}\nabla_{\perp}^{2}\eta-\hat{h}_{-}\eta-\mathbf{P}_{\perp}\cdot\mathbf{\sigma}_{\perp}\eta$ (24)
@@
Fig.(2) presents a schematic plot of helicity states of photon and its mirror counterpart - anti-photon. In general, within spin-degenerate media, the helical states of photons evolve independently. However, when considering the state of the anti-photon, a transition may occur between the helical states of the photon and the anti-photon during propagation, as shown in Equation (LABEL:29). This transition arises from the fact that for a right-handed helical state ($\xi$) the corresponding state for photons traveling in the opposite direction is left-handed. Consequently, the helical state of the anti-photon can be incorporated into the opposite helical state of the photon, the anti-photon effectively representing back-scattering. Therefore, the states $\xi$ and $\eta$ evolve independently during propagation, maintaining their respective identities within the medium.
III EQUATION OF MOTION

### v2

@@ I Introduction
Polarization refers to the orientation of an electromagnetic wave’s electric/magnetic field vector. In the context of quantum theory, it corresponds to the spin angular momentum of photons, an intrinsic degree of freedom of light encoded in the classical Maxwell theory. Adapting our understanding of electromagnetic waves to incorporate the concept of angular momentum is critical for comprehending the intricate dynamics at play. Besides the spin angular momentum (SAM) of an electromagnetic wave packet, we have appreciated another two pivotal components: intrinsic orbital angular momentum(IOAM) associated with optical vortices, which characterizes the inherent rotation of light around its axis, and extrinsic orbital angular momentum (EOAM) corresponds to the helical optical trajectory, extending beyond its inherent properties to encompass its path through space [1, 2, 3, 4, 5, 6, 7]. Interactions among these three angular momenta give rise to a diverse range of optical phenomena and can be manipulated to control various aspects of optical systems.
@@
The spin-orbit interaction (SOI) of light manifests itself through the interplay and mutual conversion between three types of angular momentum[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], a family of fundamental phenomena that arises directly from the solutions of Maxwell’s equations, particularly when dealing with wavelength-scale micro- and nanostructures. Obviously, a full treatment of the SOI phenomenon on these scales requires vector-wave mechanics. This motivation has led us to develop an extended Dirac theory for optical fields in generic media, which was found to be a non-Hermitian chiral extension of massive fermions with anomalous magnetic moments moving in an external pseudo-magnetic field [19]. In this approach, an optical medium prescribed by a symmetric dielectric tensor can be decomposed in terms of the monopole, dipole, and quadrupole components, corresponding to the spin-0 scalar mode, spin-1 vector mode, and spin-2 tensor mode, respectively. While coupling to optical fields, those virtual excitation modes ensure the conservation of total angular momentum (TAM). On the other hand, the chiral $\gamma_{5}$ extension leads to a scenario of an anomalous magnetic momentum in an equivalent magnetic field, a synthetic one in a non-spin-degenerate medium.
@@
This study aims to develop a field-theoretical representation of spin-orbit interaction (SOI) and the spin/orbit Hall effects (SHE/OHE) within the framework of optical Dirac theory, providing deeper physical insights into related optical phenomena. To this end, we extend optical Dirac theory [19] by incorporating antisymmetric components in the dielectric tensor. These components manifest either as a rotational vector in the spin-1 mode, associated with a helical optical path, or as a spin-2 tensor mode, corresponding to gyro-electric/magnetic media.
In Section II, we derive the optical Dirac equation in a helical coordinate system with a rotational dipole-vector, leading to an effective Hamiltonian that explicitly captures both SAM-EOAM and IOAM-EOAM interactions. We also provide leading-order solutions to the optical Dirac equation, incorporating SAM/IOAM-EOAM interactions. By retaining longitudinal corrections, we further examine back-scattering phenomena related to the anti-photon state. In Section III, we employ the Heisenberg equation to derive the photon motion equation, introducing characteristic corrections to light-ray trajectories, which manifest as the spin Hall effect (SHE) and orbital Hall effect (OHE). Section IV explores the evolution of SAM and IOAM in the Heisenberg picture, elucidating their role in the broader SHE/OHE family. In Section V, we analyze the interaction between helicity-independent spin and EOAM, revealing behavior analogous to that of IOAM. Finally, Section VI summarizes our findings and presents concluding remarks. The extension to gyro-electric/magnetic media follows naturally, as shown in Appendix A.
@@ II OPTICAL DIRAC EQUATION IN HELICAL COORDINATES
For light propagation in a gravitational field, the curved space background can be transformed to a linear optical medium, whose optical properties characterized by the effective dielectric tensor are fully specified by the spacetime geometries [20]. Typically, in a rotating gravitational field, a well-known optical phenomenon is the classical phase shift referred to as the Sagnac factor [21], currently explained by the dragging effect of the frame in general relativity [22], which can be understood alternatively by the coupling between the rotation and the orbital angular momentum(OAM) of photons $\sim{\bf\Omega}\cdot{\bf L}$. In addition, another phenomenon is an extra polarization-dependent phase shift resulting from the vortical dragging vector, which leads to the optical activity of rotating gravitational fields [23, 24, 25]. Basically, it can also be equivalently described by the coupling between the macroscopic rotation ${\bf\Omega}$ and the internal spin ${\bf s}$ of the photon [26, 24]. This coupling has a typical form of $\sim{\bf\Omega}\cdot{\bf s}$.
Figure 1: The local coordinate system is constructed from the Frenet frame $\mathbf{t},\mathbf{n},\mathbf{b}$, where $C$ denotes the trajectory of the ray system of light. The position vector $\mathbf{r}_{c}$ describes the ray’s location relative to the laboratory reference frame, whose origin is at point $O$. The transverse displacement vector $\mathbf{r}_{\perp}$ characterizes the deviation perpendicular to the ray direction. The blue-shaded region, spanned by the transverse displacement vectors along the ray path, defines a three-dimensional volume that can be conceptually likened to an optical fiber.
It is noted that both rotation-OAM/SAM couplings are due to spatial rotation in time, which can be attributed to the dragging vector, the space-time component ${\bf g}=\{g_{0i}\}$ in metric. Similarly, a similar rotation occurs in space for a helical optical path. As demonstrated later, this rotation around a spatial axis will give rise to a similar coupling effect, e.g., the optical SHE. Suppose light propagates along the helical path in space (see FIG. 1). Under the geometric optics approximation, an optical path can be parameterized by its optical length $s$, and let $\mathbf{r}_{\perp}=\{x_{1},x_{2}\}$ denote a transverse displacement vector around the propagating direction; we can introduce local coordinates $(s,x_{1},x_{2})$, which are related to the position vector ${\mathbf{R}}=(x,y,z)$ in the laboratory frame by
@@
In the local Frenet-Serret orthonormal tetrad defined by curvature $\kappa(s)$ and torsion $\tau(s)$, the line element is $d{\bf R}^{2}=\widetilde{\gamma}_{\alpha\beta}dx^{\alpha}dx^{\beta}$ where Greek indices run over 0, 1 and 2, and $x_{0}=s$ by convention. The metric components can be easily found [27]:
$\widetilde{\gamma}_{00}=(1-\kappa x_{1})^{2}+\tau^{2}r_{\perp}^{2},\widetilde{\gamma}_{0\alpha}=-\tau\varepsilon_{0\alpha\beta}x_{\beta},\widetilde{\gamma}_{\alpha\beta}=\delta_{\alpha\beta}$ (2)
@@
${\frac{1}{\sqrt{\widetilde{\gamma}}}}\frac{\partial}{\partial t}{(\sqrt{\widetilde{\gamma}}}D^{\alpha})=\epsilon^{\alpha\beta\rho}\nabla_{\beta}\bigr{(}\sqrt{\widetilde{\gamma}}\widetilde{\gamma}_{\rho\sigma}B^{\sigma}\bigr{)}$ (3)
${\frac{1}{\sqrt{\widetilde{\gamma}}}}\frac{\partial}{\partial t}{(\sqrt{\widetilde{\gamma}}}B^{\alpha})=\epsilon^{\alpha\beta\rho}\nabla_{\beta}\bigr{(}\sqrt{\widetilde{\gamma}}\widetilde{\gamma}_{\rho\sigma}D^{\sigma}\bigr{)}$
${\frac{1}{\sqrt{\widetilde{\gamma}}}}\nabla_{\alpha}(\sqrt{\widetilde{\gamma}}D^{\alpha})=0$
${\frac{1}{\sqrt{\widetilde{\gamma}}}}\nabla_{\alpha}(\sqrt{\widetilde{\gamma}}B^{\alpha})=0$
where $\sqrt{\widetilde{\gamma}}=|\det(\widetilde{\gamma}_{\alpha\beta})|^{1/2}$ that can be absorbed into the definition of the electromagnetic vectors ${\bf D}$ and ${\bf B}$ by this rescaling. It is noted that an electromagnetic system described by Eq.(3) in curvilinear coordinates is equivalent to an optical medium with identical permittivity and permeability tensors,
$\{\epsilon^{-1}\}_{\alpha\beta}=\{\mu^{-1}\}_{\alpha\beta}=\widetilde{\gamma}_{\alpha\beta}$ (4)
which is currently referred to as the spin-degenerate condition in photonic topological insulators. Following the same method as described in Feng & Wu [19], we project the Maxwell equation onto the helicity space spanned by ${\bf e}_{\pm}=({\bf e}_{1}\pm i{\bf e}_{2})/\sqrt{2}$ where $\{{\bf e}_{1},{\bf e}_{2}\}$ is given by Eq.(1) to define the transverse tetrad perpendicular to the wavenumber vector of photon ${\bf e}_{\bf k}={\bm{\nabla}}_{s}$. Let $\bm{D}_{\perp}{=(D_{+},D_{-})}$, ${\bm{B}}_{\perp}{=(B_{+},B_{-})}$ denote the transverse components of electromagnetic field vectors in the helicity space, we introduce two-component wavefunctions for photons in the transverse helicity space:
@@
Given the permittivity and permeability tensors Eq.(4), following the derivation procedure outlined in Appendix A, the optical Dirac equation in the helical coordinate system can be expressed as
@@
in which the first term in $\gamma_{0}$ term of r.h.s. gives the leading order of the Hamiltonian, including the conventional momentum operator $\hat{k}_{z}=-i\frac{\partial}{\partial z}$ in the propagating direction and the rotation-angular momentum coupling, ${\mathbf{\Omega}}\cdot{\mathbf{J}}$, where the rotation is specified by the torsion of the helical path, $\mathbf{\Omega}=-\tau{\bf e}_{z}$, and ${\bf J}$ is the TAM, ${\bf J}={\bf L}+{\bf\Sigma}$, here ${\bf L}=\hat{\bf x}\times\hat{\bf k}$ is the OAM operator, ${\bf\Sigma}=\bm{\sigma}\otimes\sigma_{0}$ is the SAM operator; The second term $\propto 1/k_{z}$ results from the paraxial corrections, and the explicit form of $P_{0}$ and $P_{\perp}$ are given by Eqs.(124-125) in Appendix A.
@@
$\mathbf{k}_{\perp}{{=i{\bm{\kappa}}_{\perp}}}=i\kappa_{x}\mathbf{e}_{x}+i\kappa_{y}\mathbf{e}_{y}.$ (8)
According to the Hamiltonian read from Eq.(6), a non-Hermitian scalar term $-i{\bf g}_{\perp}/k_{z}$ appears in $P_{0}$ (see Eq.(125)), where ${\bf g}_{\perp}=\frac{1}{2}{\bm{\nabla}}_{\perp}\tilde{\gamma}_{00}$. However, when coupled to the evanescent wave, this term transforms into a Hermitian wave-interaction ${\bm{\kappa}}_{g}\cdot{\bm{\kappa}}{{}_{\perp}}$ with ${\bm{\kappa}}_{g}={\bf g}_{\perp}/k_{z}$, and contributes an additional energy correction.
III TWISTED LIGHT IN THE OPTICAL DIRAC THEORY
@@
with the 2-spinors $\chi_{1}=\displaystyle{\binom{1}{0}}$ and $\chi_{2}=\displaystyle{\binom{0}{1}}$. Keeping the leading order of $O(k_{z})$, we can easily see that the eigenvalues for the positive/negative energy states $u_{\alpha}$ and $v_{\alpha}$ are $\omega^{(\pm)}=\pm k$, respectively. In Dirac’s relativistic theory of the electron, the negative energy state corresponds to its anti-particle, with opposite charge, namely, the positron. However, for the spin-1 photon, the corresponding antiparticle is itself. Obviously, it can be understood as moving toward the negative arrow of time, equivalently, a backward propagation solution.
@@
$\omega^{(+)}=\omega_{\alpha}=k_{z}\mp\tau$ (11)
@@
$\omega^{(-)}=-\omega_{\alpha}=-(k_{z}\pm\tau)$ (12)
@@
$\omega_{\alpha}$ $=k_{z}\mp(l\tau+\tau)=k_{z}\mp(l+1)\tau$ (14)
$\omega_{\alpha}$ $=k_{z}\mp(l\tau-\tau)=k_{z}\mp(l-1)\tau$
These relationships can be expressed compactly as:
$\omega_{\alpha}=k-j_{z}\tau$ (15)
with $j_{z}$ representing the component of the TAM $\mathbf{J}$ along the propagating direction. The possible values of $j_{z}$ are $j_{z}\in\{-l-1,-l+1,l-1,l+1\}$.
The second term contributes to Berry phase as [29]:
@@
$ik_{z}\frac{\partial\xi}{\partial z}=-\frac{1}{2}g_{s}\nabla_{\perp}^{2}\xi-\hat{h}_{+}\xi-\mathbf{P}_{\perp}\cdot{\bm{\sigma}}_{\perp}\xi$ (21)
in which $\mathbf{P}_{\perp}\cdot{\bm{\sigma}}_{\perp}=P_{+}\sigma_{+}+P_{-}\sigma_{-}$. Similarly, defining
@@
$ik_{z}\frac{\partial\eta}{\partial z}=-\frac{1}{2}g_{s}\nabla_{\perp}^{2}\eta-\hat{h}_{-}\eta-\mathbf{P}_{\perp}\cdot{\bm{\sigma}}_{\perp}\eta$ (23)
@@
To understand the spin precession described by the interaction term ${\bf P}_{\perp}\cdot{\bm{\sigma}}_{\perp}$ of Eq.(21) and Eq.(23), it is helpful to emphasize the following three simple key points [19]: (1) The matrices $\sigma_{\pm}$ describe spin-flip transitions. Specifically, $\sigma_{+}$ induces a transition from the spin-down to the spin-up state, corresponding to the absorption of two units of angular momentum by the photon. Conversely, $\sigma_{-}$ induces the reverse transition, from spin-up to spin-down. (2) In the cylindrical coordinates $\{\rho,\phi,z\}$, we have
$\hat{k}_{\pm}=\frac{1}{\sqrt{2}}(\hat{k}_{x}\mp i\hat{k}_{y})=-\frac{i}{\sqrt{2}}e^{\mp i\phi}\left(\partial_{\rho}\pm\frac{1}{\rho}L_{z}\right)$ (25)
For the eigenstate of z-component of OAM, $L_{z}|l\rangle=l|l\rangle$, $k_{\pm}$ acts as the ladder operators to lower/raise the z-component of orbital angular momentum (OAM) by one quantum of $\hbar$. (3) Any arbitrary dielectric tensor can be decomposed into symmetric and anti-symmetric parts. The symmetric part can further be decomposed into scalar, vector, and tensor modes, each possessing two degrees of freedom. These correspond to virtual excitation modes with spin-0 (monopole), spin-1 (dipole), and spin-2 (quadrupole), respectively. Such a spin-mode classification can be found in metric gravity theories for distinguishing distinct polarization modes of gravitational waves [33]. While for the anti-symmetric component, it can be interpreted as inducing an effective rotational behavior in the medium, leading to the optical activity observed in gyrotropic media.
With these rules in place, the structure of the interaction term $P_{\pm}\cdot\sigma_{\pm}$ becomes more transparent, explicitly ensuring the conservation of total angular momentum (TAM) in each term and elucidating the underlying mechanism for the interconversion between different forms of angular momentum. For instance, in the first SAM-OAM interaction term proportional to $k_{\pm}^{2}\sigma_{\pm}$ (see Eq.(125)), a spin-flip is accompanied by a change in IOAM of the photon. A flip from spin-down to spin-up (via $\sigma_{+}$) requires the photon to gain two units of SAM, which is compensated by a reduction of two units in OAM and vice versa for $\sigma_{-}$. While for the second term, it is of the dipole-IOAM-SAM interaction form $\sim d_{\pm}k_{\pm}\sigma_{\pm}$, where $d_{\pm}$ relies on the dielectric tensor can be read from Eq.(125). Clearly, the compensation of two units of SAM through $\sigma_{\pm}$ is achieved by one unit of IOAM via the $k_{\pm}$ operator, with the remaining unit provided by the optical medium through the absorption of a spin-1 vector mode $d_{\pm}$. In the latter, the intervention of spin-1 modes of the medium leads to the change of its angular momentum, which provides a microscope explanation of the optical torque exerted by twisted light in the context of field theory.
In addition, we notice that there are another terms $h_{+}$ in Eq.(21) for $\xi$ and $h_{-}$ in Eq.(23) for $\eta$. It is easy to find that this spin-independent scalar term can be written in a compact form,
$h_{\pm}=g_{\mp}\hat{k}_{\pm}$ (26)
where $g_{\pm}$ is the components of ${\bf g}_{\perp}$ in the helicity space, i.e. ${\bf g}_{\perp}=g_{\alpha}{\bf e}_{\alpha},\alpha=+,-$. Eq.(26) describes a dipole interaction between the IOAM of twisted light and optical medium. Specifically, it corresponds to the exchange of one unit of OAM quantum number, thereby imparting an additional optical torque on the medium.
Unlike the conventional two-spinor formalism used for spin-1/2 particles, the spinors $\xi$ and $\eta$ describe two-component states of photon–antiphoton pairs with fixed positive and negative helicities, respectively. Figure 2 illustrates the schematic representation of the helicity states of a photon and its mirror counterpart—the antiphoton. In general, within spin-degenerate media, the helical states of photons propagate independently. However, when the antiphoton state is taken into account, a transition between the helical states of the photon and antiphoton can occur during propagation, as indicated in Eq. (LABEL:29). This transition stems from the fact that a right-handed helical state (represented by $\xi$) for a forward-propagating photon corresponds to a left-handed state when the photon propagates in the opposite direction. Thus, the helicity state of the antiphoton can effectively be interpreted as the opposite helicity state of the photon under time or spatial inversion—physically corresponding to backscattering. Consequently, while the antiphoton may be viewed as a reflected photon with reversed helicity, the spinor fields $\xi$ and $\eta$ remain dynamically decoupled in the absence of such backscattering processes, evolving independently and preserving their respective identities within the medium.
Subject to the conservation of TAM, the emergence of an antiphoton must be accompanied by a corresponding change in OAM relative to its companion photon. This observation leads to an intriguing phenomenological consequence: in a spin-degenerate optical medium, a structured light beam with a Gaussian profile propagating in a given direction may generate antiphoton-like modes—namely, vortex beams propagating in the opposite direction while retaining the same helicity. Conversely, another possible scenario involves vortex beams, such as vector Laguerre–Gaussian modes with $l=1$ or $l=2$, giving rise to antiphoton modes in the form of vortex-free optical fields. These are characterized by Gaussian-like beams with a bright central intensity—effectively representing the reversal of the typical doughnut-shaped intensity profile associated with optical vortices.
IV EQUATION OF MOTION